A cylinder of mass 10 kg and radius 20 cm is free to rotate about its axis. It receives an angular impulse of 4 kg m2/s. what is the angular speed of the cylinder if the cylinder is initially at rest?
(a) 20 rad/s (b) 15 rad/s
(c) 10 rad/s (d) 5 rad/s

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Vikas TU
3 years ago
It is given that,
mass of cylinder, m = 5kg
radius of cylinder , r = 30cm = 0.3 m
angular impulse , ∆L = 3kgm²/s
here, initial angular speed is 
angular speed after 4s = 
we know,
angular impulse = change in angular momentum
3 = 
Here I is moment of inertia.
3 =1/2 mr²( -0)
6 = 5 × (0.3)² × 
 = 40/3 rad/s
now, use 
40/3 = 0 +  × 4
 = 10/3 rad/s
now angular speed after 30 sec = 0 + 10/3 × 30
= 300/3 = 100 rad/s